(1) What is the Molarity of a solution formed when 58.5g of Nacl is dissolved in 200ml of water.D=(1.00g/ml)
 
(2) Calculate the mole of fraction, Molarity and Molality NH3 in 81.3g of H20. The density of the solution is 0.928g/mol and the density of water is  1.00g/ml
(3) Calculate the Molarity of each of the following (a) 12.4g Kcl in 289.2ml solution (b) 16.4g Cacl2 in 0.614L solution (c) 48.0ml of 6.00 MH2SO4 diluted to 0.250L
 
(4) Calculate the Molality of each of the following solution (a) 2.89g of Nacl dissolved in 0.159L of water (d=1.00g/ml) (b) 1.80mol Kcl in 16.0 mole of H20 (c) 130g benzene C6H6 in 17.0g CCL4
 
(5) Calculate the Molarity of a 31% HCL solution having a density of 1.2g/ml

Grade:12th pass

1 Answers

Prince choudhary
2 years ago
Q. 1 mass of Nacl =58.5g(given) 
       Volume of water=200ml=0.2litre 
        Moler mass of NaCl = 23+35.5=58.5 
       
        So number of moles of Nacl=1 
       
        Molarity =no. Of mole of Nacl
                         /volume of water( in litre) 
                        = 1/0.2=5